OpenStudy (clamin):

1.The force of the wind blowing on a vertical surface varies jointly as the area of the surface and the square of the velocity. If a wind blowing at 50 mph exerts a force of 75 pounds on a surface of 500 ft^2, how much force will a wind of 75 mph place on a surface of 10 ft^2?

OpenStudy (irishboy123):

we can start off with \[F = k A v^2\] does that make sense? incidentally this looks a lot like Rayleigh's law for drag :-)

OpenStudy (irishboy123):

\[F_D\, =\, \tfrac12\, \rho\, u^2\, C_D\, A\] i.e. \(F\, =k \, v^2\,\, A \qquad \qquad\) ;-)) based on: https://en.wikipedia.org/wiki/Drag_equation

OpenStudy (clamin):

@IrishBoy123 what do i put for A and F??

OpenStudy (irishboy123):

\[F\, =k \, v^2\,\, A \] \[v= 50 mph \quad F= 75 lbs, \quad A= 500 ft^2 \quad \text {solve for k}\] \[v= 75 mph \quad k = \text{the same} \quad A = 10 ft^2, \quad F = ???\]

OpenStudy (irishboy123):

or \[\dfrac{F_1}{ v_1^2\,\, A_1} =k = \dfrac{F_2}{ v_2^2\,\, A_2} \] solve from there...

OpenStudy (irishboy123):

and plug in some numbers to get a feel for it.